Two balls are 8.0 meters apart and moving directly towards each other. If the first ball is moving at a speed of 2.5 ms–1 with respect to the ground and the second ball 3.5 ms–1 with respect to the ground, where will they collide? 


1
Expert's answer
2021-10-18T10:57:19-0400

"s = 8m"

"v_1 =2.5\frac{m}{s}\newline
v_2 =3.5\frac{m}{s}\newline
s_1 = v_1t\newline
s_2 = v_2t\newline"

"s = s_1+s_2"

"v_1t +v_2t = s"

"t= \frac{s}{v_1+v_2}=\frac{8}{6}\approx 1.33s"

"s_1 =1.33*2.5\approx3.33m"

"s_2 =s - s_1= 4.67m"


"\text{Answer: distance collision point :}"

"3.33m \text{ from the beginning of the movement of 1 ball}"

"4.67 m \text{ from the beginning of the movement of 2 ball}"